# Optimal Computation in Leaderless and Multi-Leader Disconnected Anonymous Dynamic Networks

### 20 Feb 2024 v5

@@ 2 Definitions and Preliminaries
We will give preliminary definitions and results, and recall some properties of history trees from [24].
Processes and networks. A dynamic network is modeled by an infinite sequence $\mathcal{G}=(G_{t})_{t\geq 1}$, where $G_{t}=(V,E_{t})$ is an undirected multigraph whose vertex set $V=\{p_{1},p_{2},\dots,p_{n}\}$ is a system of $n$ anonymous processes and $E_{t}$ is a multiset of edges representing links between processes.
@@
Each process $p_{i}$ starts with an input $\lambda(p_{i})$, which is assigned to it at round 00. It also has an internal state, which is initially determined by $\lambda(p_{i})$. At each round $t\geq 1$, every process composes a message (depending on its internal state) and broadcasts it to its neighbors in $G_{t}$ through all its incident links. [^9] In order to model wireless radio communication, it is natural to assume that each process in a dynamic network broadcasts its messages to all its neighbors (a message is received by anyone within communication range). The network’s anonymity prevents processes from specifying single destinations. By the end of round $t$, each process reads all messages coming from its neighbors and updates its internal state according to a local algorithm $\mathcal{A}$. Note that $\mathcal{A}$ is deterministic and is the same for all processes. [^10] This network model bears some similarities with Population Protocols, although there are radical differences in the way symmetry is handled. The point-to-point communication model of Population Protocols automatically breaks the symmetry between communicating agents, greatly simplifying problems such as Leader Election, Average Consensus, etc. The input of each process also includes a leader flag. The processes whose leader flag is set are called leaders (or supervisors). We will denote the number of leaders as $\ell$.
@@
Each process also returns an output at the end of each round, which is determined by its current internal state. A system is said to stabilize if the outputs of all its processes remain constant from a certain round onward; note that a process’ internal state may still change even when its output is constant. A process may also decide to explicitly terminate and no longer update its internal state. When all processes have terminated, the system is said to terminate, as well.
@@
We say that $\mathcal{A}$ computes a function $F$ if, whenever the processes are assigned inputs $\lambda(p_{1})$, $\lambda(p_{2})$, …, $\lambda(p_{n})$ and all processes execute the local algorithm $\mathcal{A}$ at every round, the system eventually stabilizes with each process $p_{i}$ giving the desired output $F(p_{i},\lambda)$. [^11] Formally, a function computed by a system of $n$ processes maps $n$-tuples of input values to $n$-tuples of output values. Writing such a function as $F(p_{i},\lambda)$ emphasizes that the output of a process may depend on all processes’ inputs, as well as on the process itself. That is, different processes may give different outputs. A stronger notion of computation requires the system to not only stabilize but also to explicitly terminate with the correct output. The (worst-case) running time of $\mathcal{A}$, as a function of $n$, is the maximum number of rounds it takes for the system to stabilize (and optionally terminate), taken across all possible dynamic networks of size $n$ and all possible input assignments.
@@
Classes of functions. Let $\mu_{\lambda}=\{(z_{1},m_{1}),(z_{2},m_{2}),\dots,(z_{k},m_{k})\}$ be the multiset of all processes’ inputs. That is, for all $1\leq i\leq k$, there are exactly $m_{i}$ processes $p_{j_{1}},p_{j_{2}},\dots,p_{j_{m_{i}}}$whose input is $z_{i}=\lambda(p_{j_{1}})=\lambda(p_{j_{2}})=\dots=\lambda(p_{j_{m_{i}}})$; note that $n=\sum_{i=1}^{k}m_{i}$. A multi-aggregate function is defined as a function $F$ of the form $F(p_{i},\lambda)=\psi(\lambda(p_{i}),\mu_{\lambda})$, i.e., such that the output of each process depends only on its own input and the multiset of all processes’ inputs.
@@
The special multi-aggregate functions $F_{C}(p_{i},\lambda)=n$ and $F_{GC}(p_{i},\lambda)=\mu_{\lambda}$ are called the Counting function and the Generalized Counting function, respectively. It is known that, if a system can compute the Generalized Counting function $F_{GC}$, then it can compute any multi-aggregate function in the same number of rounds: thus, $F_{GC}$ is complete for the class of multi-aggregate functions [24].
For any $\alpha\in\mathbb{R}^{+}$, we define $\alpha\cdot\mu_{\lambda}$ as $\{(z_{1},\alpha\cdot m_{1}),(z_{2},\alpha\cdot m_{2}),\dots,(z_{k},\alpha\cdot m_{k})\}$. We say that a multi-aggregate function $F(p_{i},\lambda)=\psi(\lambda(p_{i}),\mu_{\lambda})$ is frequency-based if $\psi(z,\mu_{\lambda})=\psi(z,\alpha\cdot\mu_{\lambda})$ for every positive integer $\alpha$ and every input $z$ (see [28]). That is, $F$ depends only on the “frequency” of each input in the system, rather than on their actual multiplicities. Notable examples include statistical functions such as mean, variance, maximum, median, mode, etc. The problem of computing the mean of all input values is called Average Consensus [6, 15, 16, 17, 26, 32, 42, 43, 45, 46, 47, 51, 54].
The frequency-based multi-aggregate function $F_{R}(p_{i},\lambda)=\frac{1}{n}\cdot\mu_{\lambda}$ is called Frequency function, and is complete for the class of frequency-based multi-aggregate functions, as shown below.
@@
If $F_{R}$ can be computed (with termination), then all frequency-based multi-aggregate functions can be computed (with termination) in the same number of rounds, as well.
@@
Suppose that a process $p_{i}$ has determined $\frac{1}{n}\cdot\mu_{\lambda}=\{(z_{1},m_{1}/n),(z_{2},m_{2}/n),\dots,(z_{k},m_{k}/n)\}$. Then it can immediately find the smallest integer $d>0$ such that $d\cdot(m_{i}/n)$ is an integer for all $1\leq i\leq k$. Note that $\frac{d}{n}\cdot\mu_{\lambda}$ is a multiset. Hence, in the same round, $p_{i}$ can compute any desired function $\psi(\lambda(p_{i}),\frac{d}{n}\cdot\mu_{\lambda})$, and thus any frequency-based multi-aggregate function, by definition. ∎
Figure 1: The first rounds of a dynamic network with $n=9$ processes and the corresponding levels of the history tree. Level $L_{t}$ consists of all nodes at distance $t+1$ from the root $r$. The multiplicities of the red multi-edges of the history tree are explicitly indicated only when greater than $1$. The letters A, B, C denote processes’ inputs; all other labels have been added for the reader’s convenience, and indicate classes of indistinguishable processes (non-trivial classes are also indicated by dashed blue lines). Note that the two processes in $b_{4}$ are still indistinguishable at the end of round $2$, although they are linked to the distinguishable processes $b_{5}$ and $b_{6}$. This is because such processes were in the same class $a_{5}$ at round $1$. The subgraph in the green blob is the view of the two processes in $b_{1}$.
History trees. History trees were introduced in [24] as a tool of investigation for anonymous dynamic networks; an example is found in Figure 1. A history tree is a representation of a dynamic network given some inputs to the processes. It is an infinite graph whose nodes are partitioned into levels $L_{t}$, with $t\geq-1$; each node in $L_{t}$ represents a class of processes that are indistinguishable at the end of round $t$ (with the exception of $L_{-1}$, which contains a single node $r$ representing all processes). The definition of distinguishability is inductive: at the end of round 00, two processes are distinguishable if and only if they have different inputs. At the end of round $t\geq 1$, two processes are distinguishable if and only if they were already distinguishable at round $t-1$ or if they have received different multisets of messages at round $t$.
Each node in level $L_{0}$ has a label indicating the input of the processes it represents. There are also two types of edges connecting nodes in adjacent levels. The black edges induce an infinite tree rooted at node $r\in L_{-1}$ which spans all nodes. The presence of a black edge $\{v,v^{\prime}\}$, with $v\in L_{t}$ and $v^{\prime}\in L_{t+1}$, indicates that the child node $v^{\prime}$ represents a subset of the processes represented by the parent node $v$. The red multi-edges represent communications between processes. The presence of a red edge $\{v,v^{\prime}\}$ with multiplicity $m$, with $v\in L_{t}$ and $v^{\prime}\in L_{t+1}$, indicates that, at round $t+1$, each process represented by $v^{\prime}$ receives $m$ (identical) messages from processes represented by $v$.
As time progresses and processes exchange messages, they are able to locally construct finite portions of the history tree. In [24], it is shown that there is a local algorithm $\mathcal{A}^{\ast}$ that allows each process to locally construct and update its own view of the history tree at every round. The view of a process $p$ at round $t\geq 0$ is the subgraph of the history tree which is spanned by all the shortest paths (using black and red edges indifferently) from the root $r$ to the node in $L_{t}$ representing $p$ (see Figure 1). As proved in [24, Theorem 3.1], the view of a process at round $t$ contains all the information that the process may be able to use at that round. This justifies the convention that all processes always execute $\mathcal{A}^{\ast}$, constructing their local view of the history tree and broadcasting (a representation of) it at every round, regardless of their task. Then, they simply compute their task-dependent outputs as a function of their respective views.
We define the anonymity of a node $v$ of the history tree as the number of processes that $v$ represents, and we denote it as $a(v)$. It follows that $\sum_{v\in L_{t}}a(v)=n$ for all $t\geq-1$, and that the anonymity of a node is equal to the sum of the anonymities of its children. Naturally, a process is not aware of the anonymities of the nodes in its view of the history tree, unless it can somehow infer them from the view’s structure itself. In fact, computing the Generalized Counting function is equivalent to determining the anonymities of all the nodes in $L_{0}$. Similarly, computing the Frequency function corresponds to determining the value $a(v)/n$ for all $v\in L_{0}$.
Computation in disconnected networks. Although the network $G_{t}$ at each individual round may be disconnected, we assume the dynamic network to be $T$-union-connected. That is, there is a dynamic disconnectivity parameter $T\geq 1$ such that the sum of any $T$ consecutive $G_{t}$’s is a connected multigraph. [^12] By definition, the sum of (multi-)graphs is obtained by adding together their adjacency matrices. Thus, for all $i\geq 1$, the multigraph $\left(V,\bigcup_{t=i}^{i+T-1}E_{t}\right)$ is connected (we remark that a union of multisets adds together the multiplicities of equal elements). [^13] Our $T$-union-connected networks should not be confused with the $T$-interval-connected networks from [36]. In those networks, the intersection (as opposed to the union) of any $T$ consecutive $E_{t}$’s induces a connected (multi)graph. In particular, a $T$-interval-connected network is connected at every round, while a $T$-union-connected network may not be, unless $T=1$. Incidentally, a network is $1$-interval-connected if and only if it is $1$-union-connected.
@@
Any non-trivial function is impossible to compute with termination unless the processes have some knowledge about $T$. (A function is “trivial” if it can be computed locally.)
@@
A function $F(p_{i},\lambda)$ is trivial if and only if it is of the form $F(p_{i},\lambda)=\psi(\lambda(p_{i}))$, i.e., the output of any process $p_{i}$ depends only on its input $\lambda(p_{i})$, and not on the inputs of other processes. Assume for a contradiction that a non-trivial function $F(p_{i},\lambda)$ is computed with termination by an algorithm $\mathcal{A}$ with no knowledge of $T$.
Since $F$ is non-trivial, there is an input value $z$ and two distinct output values $y$ and $y^{\prime}$ with the following properties. (i) If $p_{1}$ is the only process in a network (i.e., $n=1$), and $p_{1}$ is assigned input $z$ and executes $\mathcal{A}$, it terminates in $t$ rounds with output $y$. (ii) There exists a network size $\tilde{n}>1$ and an input assignment $\lambda$ with $\lambda(p_{1})=z$ such that, whenever the processes in a network of size $\tilde{n}$ are assigned input $\lambda$ and execute $\mathcal{A}$, the process $p_{1}$ eventually terminates with output $y^{\prime}\neq y$.
Let us now consider a dynamic network of $\tilde{n}$ processes, where $p_{1}$ is kept disconnected from all other processes for the first $t$ rounds (hence $T>t$). Assign input $\lambda$ to the processes, and let them execute algorithm $\mathcal{A}$. Due to property (i), since $p_{1}$ is isolated for $t$ rounds and has no knowledge of $T$, it terminates in $t$ rounds with output $y$. This contradicts property (ii), which states that $p_{1}$ should terminate with output $y^{\prime}\neq y$. ∎
@@
A function $F$ can be computed (with termination) within $f(n)$ rounds in any dynamic network with $T=1$ if and only if $F$ can be computed (with termination) within $T\cdot f(n)$ rounds in any dynamic network with $T\geq 1$, assuming that $T$ is known to all processes.
@@
Subdivide time into blocks of $T$ consecutive rounds, and consider the following algorithm. Each process collects and stores all messages it receives within a same block, and updates its state all at once at the end of the block. This reduces any $T$-union-connected network $\mathcal{G}=((V,E_{t}))_{t\geq 1}$ to a $1$-union-connected network $\mathcal{G}^{\prime}=((V,E^{\prime}_{t}))_{t\geq 1}$, where $E^{\prime}_{t}=\bigcup_{i=(t-1)T+1}^{tT}E_{i}$. Thus, if $F$ can be computed within $f(n)$ rounds in all $1$-union-connected networks (which include $\mathcal{G}^{\prime}$), then $F$ can be computed within $Tf(n)$ rounds in the original network $\mathcal{G}$ [^14] Note that this argument is correct because algorithms are required to work for all multigraphs, as opposed to simple graphs only. Indeed, since a process $p_{i}$ may receive multiple messages from the same process $p_{j}$ within a same block, the resulting network $\mathcal{G}^{\prime}$ may have multiple links between $p_{i}$ and $p_{j}$ in a same round, even if $\mathcal{G}$ does not.
@@
Conversely, consider a $1$-union-connected network $\mathcal{G}$, and construct a $T$-union-connected network $\mathcal{G}^{\prime}$ by inserting $T-1$ empty rounds (i.e., rounds with no links at all) between every two consecutive rounds of $\mathcal{G}$. Since no information circulates during the empty rounds, if $F$ cannot be computed within $f(n)$ rounds in $\mathcal{G}$, then $F$ cannot be computed within $Tf(n)$ rounds in $\mathcal{G}^{\prime}$ (recall that running times are measured in the worst case across all possible networks). ∎
@@
Relationship with the dynamic diameter. A concept closely related to the dynamic disconnectivity $T$ of a network is its dynamic diameter (or temporal diameter) $D$, which is defined as the maximum number of rounds it may take for information to travel from any process to any other process at any point in time [10, 38]. It is a simple observation that $T\leq D\leq T(n-1)$.

### 05 Apr 2024 v6

@@ 2 Definitions and Preliminaries
In Section 2.1 we define the model of computation of anonymous dynamic networks. In Section 2.2 we define two classes of functions and their respective “complete” functions, which will play an important role in the rest of the paper. Finally, in Section 2.3 we discuss $\tau$-union-connected networks and the impact of the dynamic disconnectivity $\tau$ on the running time of algorithms.
2.1 Model of Computation
Processes and networks. A dynamic network on a finite non-empty set $V=\{p_{1},p_{2},\dots,p_{n}\}$ is an infinite sequence $\mathcal{G}=(G_{t})_{t\geq 1}$, where $G_{t}=(V,E_{t})$ is an undirected multigraph, i.e., $E_{t}$ is a multiset of unordered pairs of elements of $V$. In this context, the set $V$ is called system, and its $n\geq 1$ elements are the processes. The elements of the multiset $E_{t}$ are called links; note that we allow any (finite) number of “parallel links” between two processes, as well as “self-loops”. [^11] In the dynamic networks literature, $G_{t}$ is typically assumed to be a simple graph, for at most one link between the same two processes is allowed. However, our results hold more generally for multigraphs.
@@ 2.1 Model of Computation
Input and internal states. Each process $p_{i}$ starts with an input $\lambda(p_{i})$, which is assigned to it at round 00. It also has an internal state, which is initially determined by $\lambda(p_{i})$. At each round $t\geq 1$, every process composes a message (depending on its internal state) and broadcasts it to its neighbors in $G_{t}$ through all its incident links. [^12] In order to model wireless radio communication, it is natural to assume that each process in a dynamic network broadcasts its messages to all its neighbors (a message is received by anyone within communication range). The network’s anonymity prevents processes from specifying single destinations. By the end of round $t$, each process reads all messages coming from its neighbors and updates its internal state according to a local algorithm $\mathcal{A}$. That is, $\mathcal{A}$ takes as input the current internal state of the process and the multiset of incoming messages, and returns the new internal state of the process. Note that $\mathcal{A}$ is deterministic and is the same for all processes. [^13] This network model bears some similarities with Population Protocols, although there are radical differences in the way symmetry is handled. The point-to-point communication model of Population Protocols automatically breaks the symmetry between communicating agents, greatly simplifying problems such as Leader Election, Average Consensus, etc. The input of each process also includes a leader flag. The processes whose leader flag is set are called leaders (or supervisors). We will denote the number of leaders by $\ell$.
@@
Stabilization and termination. Each process returns an output at the end of each round, which is determined by its current internal state. A system is said to stabilize if the outputs of all its processes remain constant from a certain round onward; note that a process’ internal state may still change even when its output is constant. A process may also decide to explicitly terminate and no longer update its internal state. When all processes have terminated, the system is said to terminate, as well.
@@
Computation. We say that an algorithm $\mathcal{A}$ computes a function $F$ if, whenever the processes in the system are assigned inputs $\lambda(p_{1})$, $\lambda(p_{2})$, …, $\lambda(p_{n})$ and all processes execute the local algorithm $\mathcal{A}$ at every round, the system eventually stabilizes with each process $p_{i}$ giving the desired output $F(p_{i},\lambda)$. [^14] Formally, a function computed by a system of $n$ processes maps $n$-tuples of input values to $n$-tuples of output values. Writing such a function as $F(p_{i},\lambda)$ emphasizes that the output of a process may depend on all processes’ inputs, as well as on the process itself. That is, different processes may give different outputs. A stronger notion of computation requires the system to not only stabilize but also to explicitly terminate with the correct output. The (worst-case) running time of $\mathcal{A}$, as a function of $n$, is the maximum number of rounds it takes for the system to stabilize (and optionally terminate), taken across all possible dynamic networks of $n$ processes and all possible input assignments.
2.2 Classes of Functions
@@
Multiset-based functions. Let $\mu_{\lambda}=\{(z_{1},m_{1}),(z_{2},m_{2}),\dots,(z_{k},m_{k})\}$ be the multiset of all processes’ inputs. That is, for all $1\leq i\leq k$, there are exactly $m_{i}$ processes $p_{j_{1}},p_{j_{2}},\dots,p_{j_{m_{i}}}$whose input is $z_{i}=\lambda(p_{j_{1}})=\lambda(p_{j_{2}})=\dots=\lambda(p_{j_{m_{i}}})$; note that $n=\sum_{i=1}^{k}m_{i}$. A function $F$ is multiset-based if it has the form $F(p_{i},\lambda)=\psi(\lambda(p_{i}),\mu_{\lambda})$. That is, the output of each process depends only on its own input and the multiset of all processes’ inputs. [^15] Multiset-based functions are also know as multi-aggregation or multi-aggregate functions [23, 25].
@@
The special multiset-based functions $F_{C}(p_{i},\lambda)=n$ and $F_{IM}(p_{i},\lambda)=\mu_{\lambda}$ are called the Counting function and the Input Multiset function, respectively. It is easy to see that, if a system can compute the Input Multiset function $F_{IM}$, then it can compute any multiset-based function in the same number of rounds: thus, $F_{IM}$ is complete for the class of multiset-based functions.
@@
If the Input Multiset function $F_{IM}$ can be computed (with termination), then all multiset-based functions can be computed (with termination) in the same number of rounds, as well.
@@
Once a process $p_{i}$ with input $\lambda(p_{i})$ has determined the multiset $\mu_{\lambda}$ of all processes’ inputs, it can immediately compute any desired function $\psi(\lambda(p_{i}),\mu_{\lambda})$ within the same local computation phase. ∎
We remark that Proposition 2.1 does not require a unique leader or a connected network.
Frequency-based functions. For any $\alpha\in\mathbb{R}^{+}$, we define $\alpha\cdot\mu_{\lambda}$ as $\{(z_{1},\alpha\cdot m_{1}),(z_{2},\alpha\cdot m_{2}),\dots,(z_{k},\alpha\cdot m_{k})\}$. We say that a multiset-based function $F(p_{i},\lambda)=\psi(\lambda(p_{i}),\mu_{\lambda})$ is frequency-based if $\psi(z,\mu_{\lambda})=\psi(z,\alpha\cdot\mu_{\lambda})$ for every positive integer $\alpha$ and every input $z$ (see [30]). That is, $F$ depends only on the “frequency” of each input in the system, rather than on their actual multiplicities. Notable examples of frequency-based functions include statistical functions of input values, such as mean, variance, maximum, median, mode, etc. Observe that the sum of all input values is a multiset-based function but not a frequency-based function.
The problem of computing the mean of all input values is called Average Consensus  [7, 16, 26, 34, 44, 45, 48, 49, 50, 55, 58]. The frequency-based function $F_{IF}(p_{i},\lambda)=\frac{1}{n}\cdot\mu_{\lambda}$ is called Input Frequency function, and is complete for the class of frequency-based functions.
@@
If the Input Frequency function $F_{IF}$ can be computed (with termination), then all frequency-based functions can be computed (with termination) in the same number of rounds, as well.
@@
Suppose that a process $p_{i}$ has determined $\frac{1}{n}\cdot\mu_{\lambda}=\{(z_{1},m_{1}/n),(z_{2},m_{2}/n),\dots,(z_{k},m_{k}/n)\}$. Then it can immediately find the smallest integer $d>0$ such that $d\cdot(m_{i}/n)$ is an integer for all $1\leq i\leq k$. Note that $\frac{d}{n}\cdot\mu_{\lambda}$ is a multiset. Hence, in the same round, $p_{i}$ can compute any desired function $\psi(\lambda(p_{i}),\frac{d}{n}\cdot\mu_{\lambda})$, and thus any frequency-based function, by definition. ∎
2.3 Disconnected Networks
Although the network $G_{t}$ at each individual round may be disconnected, in this paper we assume dynamic networks to be $\tau$-union-connected. That is, there is a (smallest) parameter $\tau\geq 1$, called dynamic disconnectivity, such that the sum of any $\tau$ consecutive $G_{t}$’s is a connected multigraph. [^16] By definition, the sum of (multi)graphs is obtained by adding together their adjacency matrices. Thus, for all $i\geq 1$, the multigraph $\left(V,\bigcup_{t=i}^{i+\tau-1}E_{t}\right)$ is connected (we remark that a union of multisets adds together the multiplicities of equal elements). [^17] Our $\tau$-union-connected networks should not be confused with the $\tau$-interval-connected networks from [38]. In those networks, the intersection (as opposed to the union) of any $\tau$ consecutive $E_{t}$’s induces a connected (multi)graph. In particular, a $\tau$-interval-connected network is connected at every round, while a $\tau$-union-connected network may not be, unless $\tau=1$. Incidentally, a network is 1-interval-connected if and only if it is 1-union-connected.
We will prove that terminating computations are impossible without some a-priori knowledge about $\tau$. We say that a function $F(p_{i},\lambda)$ is trivial if and only if it is of the form $F(p_{i},\lambda)=\psi(\lambda(p_{i}))$. That is, the output of any process $p_{i}$ depends only on its own input $\lambda(p_{i})$ and not on the inputs of other processes; therefore, $F$ can be computed “locally” and does not require communications between processes.
@@ 2.3 Disconnected Networks
Any non-trivial function is impossible to compute with termination unless the processes have some knowledge about $\tau$.
@@
Assume for a contradiction that a non-trivial function $F(p_{i},\lambda)$ is computed with termination by an algorithm $\mathcal{A}$ with no knowledge of $\tau$. Since $F$ is non-trivial, there is an input value $z$ and two distinct output values $y$ and $y^{\prime}$ with the following properties. (i) If $p_{1}$ is the only process in a network (i.e., $n=1$), and $p_{1}$ is assigned input $z$ and executes $\mathcal{A}$, it terminates in $t$ rounds with output $y$. (ii) There exists a network size $\tilde{n}>1$ and an input assignment $\lambda$ with $\lambda(p_{1})=z$ such that, whenever the processes in a network of size $\tilde{n}$ are assigned input $\lambda$ and execute $\mathcal{A}$, the process $p_{1}$ eventually terminates with output $y^{\prime}\neq y$.
Let us now consider a dynamic network of $\tilde{n}$ processes, where $p_{1}$ is kept disconnected from all other processes for the first $t$ rounds (hence $\tau>t$). Assign input $\lambda$ to the processes, and let them execute algorithm $\mathcal{A}$. Due to property (i), since $p_{1}$ is isolated for $t$ rounds and has no knowledge of $\tau$, it terminates in $t$ rounds with output $y$. This contradicts property (ii), which states that $p_{1}$ should terminate with output $y^{\prime}\neq y$. ∎
The following result will be used repeatedly to reduce computations on $\tau$-union connected networks to computations on 1-union-connected networks.
Proposition 2.4.
A function $F$ can be computed (with termination) within $f(n)$ rounds in any dynamic network with $\tau=1$ if and only if $F$ can be computed (with termination) within $\tau\cdot f(n)$ rounds in any dynamic network with $\tau\geq 1$, assuming that $\tau$ is known to all processes.
Proof.
@@
Subdivide time into blocks of $\tau$ consecutive rounds, and consider the following algorithm. Each process collects and stores all messages it receives within a same block, and updates its state all at once at the end of the block. This reduces any $\tau$-union-connected network $\mathcal{G}=((V,E_{t}))_{t\geq 1}$ to a 1-union-connected network $\mathcal{G}^{\prime}=((V,E^{\prime}_{t}))_{t\geq 1}$, where $E^{\prime}_{t}=\bigcup_{i=(t-1)\tau+1}^{t\tau}E_{i}$. Thus, if $F$ can be computed within $f(n)$ rounds in all 1-union-connected networks (which include $\mathcal{G}^{\prime}$), then $F$ can be computed within $\tau f(n)$ rounds in the original network $\mathcal{G}$ [^18] Note that this argument is correct because algorithms are required to work for all multigraphs, as opposed to simple graphs only. Indeed, since a process $p_{i}$ may receive multiple messages from the same process $p_{j}$ within a same block, the resulting network $\mathcal{G}^{\prime}$ may have multiple links between $p_{i}$ and $p_{j}$ in a same round, even if $\mathcal{G}$ does not.
@@
Conversely, consider a 1-union-connected network $\mathcal{G}$, and construct a $\tau$-union-connected network $\mathcal{G}^{\prime}$ by inserting $\tau-1$ empty rounds (i.e., rounds with no links at all) between every two consecutive rounds of $\mathcal{G}$. Since no information circulates during the empty rounds, if $F$ cannot be computed within $f(n)$ rounds in $\mathcal{G}$, then $F$ cannot be computed within $\tau f(n)$ rounds in $\mathcal{G}^{\prime}$ (recall that running times are measured in the worst case across all possible networks). ∎
@@
Relationship between dynamic disconnectivity and the dynamic diameter. A concept closely related to the dynamic disconnectivity $\tau$ of a network is its dynamic diameter (or temporal diameter) $d$, which is defined as the maximum number of rounds it may take for information to travel from any process to any other process at any point in time [11, 40]. It is a simple observation that $\tau\leq d\leq\tau(n-1)$.